Richard's average driving speed is 23 km per hour faster than Kirk's. In the same length of time it takes Richard to drive 651 km, Kirk drives on 490 km. What is Richard's average speed?
Answer by checkley77(12569)

651/(X+23)=490/X
651X=490(X+23)
651X=490X+11,270
651X-490X=11,270
161X=11,270
X=11,270/161
X=70 KMH FOR kIRK'S SPEED
70+23=93 KMH FOR rICHARDS SPEED.
PROOF:
651/93=490/70
7=7